### v1

# Tightness for interlacing geometric random walk bridges

Abstract.
We investigate a class of line ensembles whose local structure is described by independent geometric random walk bridges, which have been conditioned to interlace with each other. The latter arise naturally in the context Schur processes, including their versions in a half-space and a finite interval with free or periodic boundary conditions. We show that under one-point tightness of the curves, these line ensembles are tight and any subsequential limit satisfies the Brownian Gibbs property. As an application of our tightness results, we show that sequences of spiked Schur processes, that were recently considered in [Dim24], converge uniformly over compact sets to the Airy wanderer line ensembles.
1. Introduction and main results
1.1. Introduction
A Gibbsian line ensemble is a collection of independent labeled random walks or Brownian motions whose joint law is reweighed by a Radon-Nikodym derivative proportional to the exponential of the sum of local interaction energies between consecutively labeled curves. A simple example of a Gibbsian line ensemble is a collection of independent random walks, conditioned not to cross each other (e.g. level lines of random lozenge or domino tilings). In this case the local energy is infinity or zero depending on whether the crossing occurs or does not. A considerably more involved example (which arises under edge scaling of certain avoiding random walks) is given by the (parabolic) Airy line ensemble, which is a collection of random continuous functions that are globally parabolic and whose local distribution is described by independent Brownian bridges, conditioned on avoiding each other (this is called the Brownian Gibbs property). The Airy line ensemble was introduced in [CH14] and it forms the foundation of the entire Kardar-Parisi-Zhang (KPZ) universality class through its role in the construction of the Airy sheet in [DOV22].
Over the last several years, there has been a substantial interest in finding general conditions that ensure that a sequence of Gibbsian line ensembles is tight, see e.g. [CH14, CH16, DFF+21, DW21b, DNV23, Ser23, Wu23a, Wu23b]. In the earliest of these works, [CH14], establishing tightness was an important ingredient in the construction of the Airy line ensemble, namely in showing that this processes has a continuous version. In some of the subsequent works establishing tightness can be seen as an important step towards showing that a sequence of Gibbsian line ensembles converge uniformly over compact sets to the Airy line ensemble, which should be thought of as a stronger notion of convergence than in the finite-dimensional sense. As mentioned in the introduction of [DNV23], establishing this stronger notion of convergence is not merely cosmetic, but a crucial ingredient in showing that various models converge to the Airy sheet. The earliest result in this direction is [DOV22], see also [ACH24] and the references therein.
Showing that a sequence of line ensembles is tight ensures the existence of subsequential limits, and to prove uniform over compacts convergence one has to show that all of these limits are the same, i.e. establish uniqueness. One way to accomplish this is to prove finite-dimensional convergence, which is hard for many models due to the lack of exact formulas. Because of this reason, a general scheme for showing uniqueness of subsequential limits has been to establish that each subsequential limit has to satisfy enough properties that narrows down the list of possible candidates to exactly one. For example, in many of the aforementioned works the authors were able to show that any subsequential limit satisfies the Brownian Gibbs property, enjoyed by the Airy line ensemble. As shown in [DM21], any line ensemble possessing this property is uniquely characterized by the law of its lowest-indexed curve, and this characterization was instrumental, together with [QS23], in establishing the convergence of the KPZ line ensemble to the Airy line ensemble in [Wu23a]. More recently, [AH23] showed that the Airy line ensemble is essentially the unique line ensemble that possesses the Brownian Gibbs property and whose top curve is globally parabolic. The latter strong characterization of the Airy line ensemble is what enabled the results in [ACH24].
The recent characterization results for Brownian Gibbsian line ensembles have placed even higher emphasis on finding general conditions that ensure that a sequence of line ensembles is tight and that any subsequential limit satisfies the Brownian Gibbs property, and the goal of this paper is to find such conditions for interlacing geometric random walk bridges.
In the present paper we deal with discrete line ensembles $\mathfrak{L}=\{L_{i}\}_{i=1}^{\infty}$, where $L_{i}:\mathbb{Z}\rightarrow\mathbb{Z}$ satisfy
(1.1) $L_{i}(s)\leq L_{i}(s+1)\mbox{ for all }s\in\mathbb{Z}.$
In words, equation (1.1) says that one can think of $L_{i}$ as the trajectories of geometric random walkers. In addition, our ensemble $\mathfrak{L}$ satisfies the property that for each $s\in\mathbb{Z}$ the vectors $\mathfrak{L}(s)=(L_{1}(s),L_{2}(s),\dots)$ and $\mathfrak{L}(s+1)=(L_{1}(s+1),L_{2}(s+1),\dots)$ interlace in the sense that
(1.2) $L_{1}(s+1)\geq L_{1}(s)\geq L_{2}(s+1)\geq L_{2}(s)\geq L_{3}(s+1)\geq L_{3}(s)\geq\cdots,$
and locally the distribution of $\mathfrak{L}$ is uniform subject to the interlacing inequalities in (1.2). The last statement will be made precise in Definition 2.6, and we refer to it as the interlacing Gibbs property. In words, $\mathfrak{L}$ satisfying the interlacing Gibbs property means that its law is locally described by independent geometric random walkers, which have been conditioned to interlace. By linearly interpolating the points $(s,L_{i}(s))$ we can view $\mathfrak{L}$ as a continuous line ensemble, see Figure 1.
[…]
Figure 1. The figure depicts the top three curves in $\mathfrak{L}$, satisfying (1.1) and (1.2).
The main result of the paper, Theorem 1.1 (which is a special case of the more general Theorem 5.1 in the main text), gives general conditions under which a sequence of line ensembles that satisfy the interlacing Gibbs property is tight (under an appropriate diffuse scaling). We furthermore show that any subsequential limit satisfies the Brownian Gibbs property. I.e. in a diffuse scaling the local geometric random walk structure of the paths becomes that of Brownian motions (or rather bridges), and the local interlacing condition becomes a locally avoiding one.
Part of the interest in ensembles that satisfy the interlacing Gibbs property comes from the fact that they arise in the Schur processes, introduced by Okounkov and Reshetikhin in [OR03]. In particular, in Section 6 we apply Theorem 1.1 to a sequence of spiked Schur processes, that were recently considered in [Dim24], and show that the latter converge uniformly over compact sets to the Airy wanderer line ensembles – see Proposition 6.2. We mention that the Airy wanderer line ensembles were constructed for general parameters in [Dim24] as weak limits of such ensembles for special parameters from [AFVM10, CH14]. By combining [Dim24, Theorem 1.8] and Proposition 6.2 one obtains an alternative construction of these ensembles, directly as weak limits of Schur processes, which is independent of the results in [AFVM10, CH14]. In addition, the more general Theorem 5.1 in the main text allows one to deal with line ensembles defined on a half-infinite or even a finite interval, which makes it applicable to a large class of models including the half-space and free boundary Schur processes from [BBCS18, BBNV18], and the periodic Schur processes from [BB19, Bor07]. In the upcoming work [DY24] we use the results of this paper to construct half-space analogues of the Airy line ensemble as weak limits of the half-space Schur processes.
1.2. Main result
As mentioned in Section 1.1 the main technical result of the paper is Theorem 5.1, which we forego stating until the main body of text after we have introduced some relevant notation in Section 2. In this section we state a special case of that result as Theorem 1.1, which provides a good idea of the type of assumptions we make and consequences we can deduce, while being easier to formulate.
A partition is a sequence $\lambda=(\lambda_{1},\lambda_{2},\dots)$ of non-negative integers such that $\lambda_{1}\geq\lambda_{2}\geq\cdots$ and all but finitely many terms are zero. We denote the set of all partitions by $\mathbb{Y}$. We say that $\lambda,\mu\in\mathbb{Y}$ interlace, denoted by $\lambda\succeq\mu$ or $\mu\preceq\lambda$, if $\lambda_{1}\geq\mu_{1}\geq\lambda_{2}\geq\mu_{2}\geq\cdots$.
We assume that $\mathfrak{L}^{N}$ is a sequence of random elements in $\mathbb{Y}^{\mathbb{Z}}$, where we write $\mathfrak{L}^{N}=\{\mathfrak{L}^{N}(s):s\in\mathbb{Z}\}$ and $\mathfrak{L}^{N}(s)=(L^{N}_{1}(s),L^{N}_{2}(s),\dots)\in\mathbb{Y}$. We make the following assumptions about this sequence.
Assumption 1. (One-point tightness) There exists $p\in(0,\infty)$, $\sigma=\sqrt{p(1+p)}$ and sequences $d_{N}\in(0,\infty),C_{N}\in\mathbb{R}$ with $\lim_{N}d_{N}=\infty$ such that for each $t\in\mathbb{R}$ and $i\in\mathbb{N}$ the sequence of random variables $\sigma^{-1}d_{N}^{-1/2}(L_{i}(\lfloor td_{N}\rfloor)-ptd_{N}-C_{N})$ is tight.
Assumption 2. (Interlacing Gibbs property) There are sequences of integers $\hat{A}_{N},\hat{B}_{N}\in\mathbb{Z}$ with $\hat{A}_{N}\leq\hat{B}_{N}$, such that $\lim_{N}d_{N}^{-1}\hat{A}_{N}=-\infty$, $\lim_{N}d_{N}^{-1}\hat{B}_{N}=\infty$, and for all $\lambda^{\hat{A}_{N}},\lambda^{\hat{A}_{N}+1},\dots,\lambda^{\hat{B}_{N}}\in\mathbb{Y}$
$\begin{split}&\mathbb{P}\left(\mathfrak{L}^{N}(\hat{A}_{N})=\lambda^{\hat{A}_{N}},\mathfrak{L}^{N}(\hat{A}_{N}+1)=\lambda^{\hat{A}_{N}+1},\dots,\mathfrak{L}^{N}(\hat{B}_{N})=\lambda^{\hat{B}_{N}}\right)\\ &=\mathbb{P}\left(\mathfrak{L}^{N}(\hat{A}_{N})=\lambda^{\hat{A}_{N}},\mathfrak{L}^{N}(\hat{B}_{N})=\lambda^{\hat{B}_{N}}\right)\cdot{\bf 1}\{\lambda^{\hat{A}_{N}}\preceq\lambda^{\hat{A}_{N}+1}\preceq\cdots\preceq\lambda^{\hat{B}_{N}}\}.\end{split}$
In plain words, the last equation states that the conditional law of $\{\mathfrak{L}^{N}(s):s=\hat{A}_{N}+1,\dots,\hat{B}_{N}-1\}$, given that $\mathfrak{L}^{N}(\hat{A}_{N})=\lambda^{\hat{A}_{N}},\mathfrak{L}^{N}(\hat{B}_{N})=\lambda^{\hat{B}_{N}}$, is uniform on $\mathbb{Y}^{\hat{B}_{N}-\hat{A}_{N}-1}$ subject to interlacing.
We can now formulate our main result about $\mathfrak{L}^{N}$.
Theorem 1.1.
Suppose that $\mathfrak{L}^{N}$ is a sequence of random elements in $\mathbb{Y}^{\mathbb{Z}}$ that satisfies Assumptions 1 and 2 above. Define the sequence $\mathcal{L}^{N}=\{\mathcal{L}_{i}^{N}\}_{i=1}^{\infty}\in C(\mathbb{N}\times\mathbb{R})$ by
$\mathcal{L}^{N}_{i}(t)=\sigma^{-1}d_{N}^{-1/2}\cdot\left(L_{i}^{N}(td_{N})-ptd_{N}-C_{N}\right),$
where $L_{i}^{N}(s)$ is defined for non-integer $s$ by linear interpolation. Then, $\mathcal{L}^{N}$ is a tight sequence of random elements in $C(\mathbb{N}\times\mathbb{R})$. Moreover, any subsequential limit satisfies the Brownian Gibbs property from [CH14], see also [DM21, Definition 2.5].
Remark 1.2.
We mention that $C(\mathbb{N}\times\mathbb{R})$ is the space of continuous functions $f:\mathbb{N}\times\mathbb{R}\rightarrow\mathbb{R}$ with the topology of uniform convergence over compact sets, see Section 2.1.
Remark 1.3.
In plain words Theorem 1.1 says that in the presence of the interlacing Gibbs property (Assumption 2), one point tightness for a sequence of line ensembles (Assumption 1) implies tightness of the sequence as random elements in $C(\mathbb{N}\times\mathbb{R})$. Moreover, the interlacing Gibbs property in the limit becomes the Brownian Gibbs property.
The proof of Theorem 1.1 is given right after Theorem 5.1 in Section 5, and is an easy corollary of the latter statement. Specifically, Theorem 5.1 generalizes Theorem 1.1 by allowing $\mathfrak{L}^{N}$ to possibly have finitely many curves, and by allowing $d_{N}^{-1}[\hat{A}_{N},\hat{B}_{N}]$ to approximate an arbitrary interval $\Lambda\subseteq\mathbb{R}$. The ability to deal with general $\Lambda$ (and not just $\Lambda=\mathbb{R}$) is important for applications to the half-space Schur processes in [BBCS18, BBNV18], where $\Lambda=(0,\infty)$, and the periodic Schur processes from [BB19, Bor07], where $\Lambda=(0,L)$ for some $L>0$.
Theorem 1.1, and more generally Theorem 5.1, is similar to [CH14, Theorem 3.8], [DFF+21, Theorem 1.1], and [Ser23, Theorem 1.1]. In [DFF+21] the authors dealt with avoiding Bernoulli random walk bridges, and [Ser23] extended the argument to a class of avoiding random walkers with general jump distributions, while [CH14] worked directly with avoiding Brownian line ensembles. One of the differences between our setup and [DFF+21, Ser23] is that our walkers are interlacing rather than avoiding – this is quite mild and does not affect the problem substantially. A more serious difference is that in [DFF+21, Ser23] the authors only worked on $\mathbb{R}$, assumed that the top curves $\mathcal{L}_{1}^{N}$ have one-point tight marginals, which are globally parabolic. This global parabolicity is what allowed them to conclude that all the other curves also need to have tight one-point marginals. In Theorem 5.1 we want to deal with possibly finite intervals $\Lambda$, for which one needs more than just information about the top curve. Specifically, it is possible for all curves except the first ones in our sequence of ensembles to dip to $\infty$- ∞, and the top curves to converge to a Brownian bridge on $\Lambda$. Consequently, one-point tightness of the top curves does not prevent other curves from escaping to $\infty$- ∞. This is why we require one-point tightness of all the curves and not just the top one. In this regard our assumptions are closer to those in [CH14].
One thing that we do differently from [CH14] is that we do not assume that the curves $\mathcal{L}^{N}_{i}$ separate from each other at a fixed time $t$ with high probability, see Hypothesis H3 in [CH14, Definition 3.3]. Instead, we show that this likely separation of the curves occurs by itself under only the one-point tightness assumption. The separation of the curves is an important ingredient in establishing tightness for $\mathcal{L}^{N}$, since it allows one to make a good comparison between interlacing random walk bridges and free (i.e. not conditioned on interlacing) random walk bridges. In particular, one can show that any event that is unlikely for the free bridges is also unlikely for the interlacing ones. The way this is used in our proofs is as follows. To show tightness, we need to show that for each $i$, the random curves $\mathcal{L}^{N}_{i}$ likely have well-behaved moduli of continuity. Being able to compare interlacing with free random walk bridges, allows us to reduce this statement to showing that a sequence of free geometric random walk bridges have well-behaved moduli of continuity. The latter statement can then be established by comparing a geometric random walk bridge with a Brownian bridge using the strong coupling from [DW21a], which we use in establishing Proposition 2.11. We mention that our arguments do not rely on the full strength of the results in [DW21a], and even a mild comparison between our random walk bridges and Brownian bridges suffices for us.
The likely curve separation was established in [DFF+21] through detailed (and rather involved) exact computations for avoiding Bernoulli random walkers, coming from Schur symmetric functions. In [Ser23] this separation was established by a careful comparison with the Bernoulli case, see [Ser23, Lemma 3.12]. Our approach is considerably simpler and free of exact computations. Specifically, it only relies on a mild comparison of our random walk bridges with Brownian bridges, and the fact that we can monotonically couple interlacing random walk bridges in their boundary. Between the two, the monotone coupling, see Lemma 2.10 for a precise statement, is the more restrictive but it is known to occur in many contexts including the aforementioned works [CH14, DFF+21, Ser23]. The crux of the separation argument is in the proof of Lemma 4.1, and the well-behaved nature of the modulus of continuity under separation is Lemma 4.3. We believe that the latter two statements and their proofs can be readily generalized to a large class of line ensembles, which satisfy the monotone coupling. This would remove the hypothesis of likely curve separation (which can be quite challenging to verify for many models), and also eliminate the need for exact computations.
1.3. Outline
In Section 2 we introduce the general class of models we study, called geometric line ensembles, and their local interlacing structure, called the interlacing Gibbs property. The section contains most of the core notation for the rest of the paper, as well as the monotone coupling, see Lemma 2.10; the comparison to Brownian bridges, see Proposition 2.11; and the statement that under diffuse scaling interlacing geometric random walk bridges converge to avoiding Brownian bridges, see Lemma 2.14. The proofs of Lemmas 2.10 and 2.14 are given in Appendix A. Section 3 provides several technical estimates for a single geometric random walk bridge, which is either free, or conditioned to stay above a well-behaved lower boundary. The proofs in Section 3 rely mostly on the monotone coupling in Lemma 2.10 and the bridge comparison in Proposition 2.11. Section 4 contains the critical Lemma 4.1, which ensures curve separation, and Lemma 4.3, which provides modulus of continuity estimates under the assumption of curve separation. Section 5 contains the most general version of our results, see Theorem 5.1. In the same section we deduce Theorem 1.1 from Theorem 5.1 and provide a proof of the latter. In Section 6 we apply Theorem 1.1 to sequences of line ensembles that arises in spiked Schur processes, and show that the latter converge to the Airy wanderer line ensembles, see Proposition 6.2.
Acknowledgments
The author would like to thank Alexei Borodin and Ivan Corwin for many inspiring conversations. This work has been partially supported by the NSF grant DMS:2230262.
2. Basic definitions and results
In this section we introduce the notions of a geometric line ensemble and the interlacing Gibbs property. Some of our notation is based on [DFF+21, Section 2], which in turn goes back to [CH14]. We also state several basic properties for these ensembles, including a monotone coupling lemma and a strong coupling of such ensembles with a single curve to Brownian bridges of appropriate variance.
2.1. Geometric line ensembles
Given two integers $a\leq b$, we let $\llbracket a,b\rrbracket$ denote the set $\{a,a+1,\dots,b\}$. We also set $\llbracket a,b\rrbracket=\emptyset$ when $a>b$, $\llbracket a,\infty\rrbracket=\{a,a+1,a+2,\dots\}$, $\llbracket-\infty,b\rrbracket=\{b,b-1,b-2,\dots\}$ and $\llbracket-\infty,\infty\rrbracket=\mathbb{Z}$. Given an interval $\Lambda\subset\mathbb{R}$, we endow it with the subspace topology of the usual topology on $\mathbb{R}$. We let $(C(\Lambda),\mathcal{C})$ denote the space of continuous functions $f:\Lambda\rightarrow\mathbb{R}$ with the topology of uniform convergence over compacts, see [Mun03, Chapter 7, Section 46], and Borel $\sigma$-algebra $\mathcal{C}$. Given a set $\Sigma\subset\mathbb{Z}$, we endow it with the discrete topology and denote by $\Sigma\times\Lambda$ the set of all pairs $(i,x)$ with $i\in\Sigma$ and $x\in\Lambda$ with the product topology. We also denote by $\left(C(\Sigma\times\Lambda),\mathcal{C}_{\Sigma}\right)$ the space of real-valued continuous functions on $\Sigma\times\Lambda$ with the topology of uniform convergence over compact sets and Borel $\sigma$-algebra $\mathcal{C}_{\Sigma}$. We typically take $\Sigma=\llbracket 1,N\rrbracket$ with $N\in\mathbb{N}\cup\{\infty\}$.
The following defines the notion of a line ensemble.
Definition 2.1.
Let $\Sigma\subseteq\mathbb{Z}$ and $\Lambda\subseteq\mathbb{R}$ be an interval. A $\Sigma$-indexed line ensemble $\mathcal{L}$ is a random variable defined on a probability space $(\Omega,\mathcal{F},\mathbb{P})$ that takes values in $\left(C(\Sigma\times\Lambda),\mathcal{C}_{\Sigma}\right)$. Intuitively, $\mathcal{L}$ is a collection of random continuous curves (sometimes referred to as lines), indexed by $\Sigma$, each of which maps $\Lambda$ in $\mathbb{R}$. We will often slightly abuse notation and write $\mathcal{L}:\Sigma\times\Lambda\rightarrow\mathbb{R}$, even though it is not $\mathcal{L}$ which is such a function, but $\mathcal{L}(\omega)$ for every $\omega\in\Omega$. For $i\in\Sigma$ we write $\mathcal{L}_{i}(\omega)=(\mathcal{L}(\omega))(i,\cdot)$ for the curve of index $i$ and note that the latter is a map $\mathcal{L}_{i}:\Omega\rightarrow C(\Lambda)$, which is $(\mathcal{C},\mathcal{F})-$measurable. If $a,b\in\Lambda$ satisfy $a\leq b$ we let $\mathcal{L}_{i}[a,b]$ denote the restriction of $\mathcal{L}_{i}$ to $[a,b]$.
Remark 2.2.
As shown in [DFF+21, Lemma 2.2], we have that $C(\Sigma\times\Lambda)$ is a Polish space, and so a line ensemble $\mathcal{L}$ is just a random element in $C(\Sigma\times\Lambda)$ in the sense of [Bil99, Section 3].
Definition 2.3.
Let $\Sigma\subseteq\mathbb{Z}$ and $\llbracket T_{0},T_{1}\rrbracket$ be a non-empty integer interval in $\mathbb{Z}$. Consider the set $Y$ of functions $f:\Sigma\times\llbracket T_{0},T_{1}\rrbracket\rightarrow\mathbb{Z}$ such that $f(i,j+1)-f(i,j)\in\mathbb{Z}_{\geq 0}$ when $i\in\Sigma$ and $j\in\llbracket T_{0},T_{1}-1\rrbracket$ and let $\mathcal{D}$ denote the discrete $\sigma$-algebra on $Y$. We call a function $g:\llbracket T_{0},T_{1}\rrbracket\rightarrow\mathbb{Z}$ such that $f(j+1)-f(j)\in\mathbb{Z}_{\geq 0}$ when $j\in\llbracket T_{0},T_{1}-1\rrbracket$ an increasing path and elements in $Y$ collections of increasing paths. A $\Sigma$-indexed geometric line ensemble $\mathfrak{L}$ on $\llbracket T_{0},T_{1}\rrbracket$ is a random variable defined on a probability space $(\Omega,\mathcal{B},\mathbb{P})$, taking values in $Y$ such that $\mathfrak{L}$ is a $(\mathcal{B},\mathcal{D})$-measurable function. Unless otherwise specified, we will assume that $T_{0}\leq T_{1}$ are both integers, although the above definition makes sense if $T_{0}=-\infty$, or $T_{1}=\infty$, or both.
Remark 2.4.
The condition $f(i,j+1)-f(i,j)\in\mathbb{Z}_{\geq 0}$ when $i\in\Sigma$ and $j\in\llbracket T_{0},T_{1}-1\rrbracket$ essentially means that for each $i\in\Sigma$ the function $f(i,\cdot)$ can be thought of as the trajectory of a geometric random walk from time $T_{0}$ to time $T_{1}$. Here, and throughout the paper a geometric random variable $X$ with parameter $q\in(0,1)$ has probability mass function $\mathbb{P}(X=k)=(1-q)q^{k}$ for $k\in\mathbb{Z}_{\geq 0}$.
We think of geometric line ensembles as collections of random increasing paths on the integer lattice, indexed by $\Sigma$. Observe that one can view an increasing path $L$ on $\llbracket T_{0},T_{1}\rrbracket$ as a continuous curve by linearly interpolating the points $(i,L(i))$, see Figure 1. This allows us to define $(\mathfrak{L}(\omega))(i,s)$ for non-integer $s\in[T_{0},T_{1}]$ and to view geometric line ensembles as line ensembles in the sense of Definition 2.1. In particular, we can think of $\mathfrak{L}$ as a random element in $C(\Sigma\times\Lambda)$ with $\Lambda=[T_{0},T_{1}]$.
We will often slightly abuse notation and write $\mathfrak{L}:\Sigma\times\llbracket T_{0},T_{1}\rrbracket\rightarrow\mathbb{Z}$, even though it is not $\mathfrak{L}$ which is such a function, but rather $\mathfrak{L}(\omega)$ for each $\omega\in\Omega$. Furthermore we write $L_{i}=(\mathfrak{L}(\omega))(i,\cdot)$ for the index $i\in\Sigma$ path. If $L$ is an increasing path on $\llbracket T_{0},T_{1}\rrbracket$ and $a,b\in\llbracket T_{0},T_{1}\rrbracket$ satisfy $a\leq b$ we let $L\llbracket a,b\rrbracket$ denote the restriction of $L$ to $\llbracket a,b\rrbracket$.
Let $t_{i},z_{i}\in\mathbb{Z}$ for $i=1,2$ be given such that $t_{1}\leq t_{2}$ and $z_{1}\leq z_{2}$. We denote by $\Omega(t_{1},t_{2},z_{1},z_{2})$ the collection of increasing paths that start from $(t_{1},z_{1})$ and end at $(t_{2},z_{2})$, by $\mathbb{P}_{\operatorname{Geom}}^{t_{1},t_{2},z_{1},z_{2}}$ the uniform distribution on $\Omega(t_{1},t_{2},z_{1},z_{2})$ and write $\mathbb{E}^{t_{1},t_{2},z_{1},z_{2}}_{\operatorname{Geom}}$ for the expectation with respect to this measure. One thinks of the distribution $\mathbb{P}_{\operatorname{Geom}}^{t_{1},t_{2},z_{1},z_{2}}$ as the law of a simple random walk with i.i.d. geometric increments with parameter $q\in(0,1)$ that starts from $z_{1}$ at time $t_{1}$ and is conditioned to end in $z_{2}$ at time $t_{2}$ – this interpretation does not depend on the choice of $q\in(0,1)$. Notice that by our assumptions on the parameters the state space $\Omega(t_{1},t_{2},z_{1},z_{2})$ is non-empty.
Given $k\in\mathbb{N}$, $T_{0},T_{1}\in\mathbb{Z}$ with $T_{0}<T_{1}$ and $\vec{x},\vec{y}\in\mathbb{Z}^{k}$ we let $\mathbb{P}^{T_{0},T_{1},\vec{x},\vec{y}}_{\operatorname{Geom}}$ denote the law of $k$ independent geometric bridges $\{B_{i}:\llbracket T_{0},T_{1}\rrbracket\rightarrow\mathbb{Z}\}_{i=1}^{k}$ from $B_{i}(T_{0})=x_{i}$ to $B_{i}(T_{1})=y_{i}$. Equivalently, this is the law of $k$ independent random increasing paths $B_{i}\in\Omega(T_{0},T_{1},x_{i},y_{i})$ for $i\in\llbracket 1,k\rrbracket$ that are uniformly distributed or just the uniform measure on
$\Omega_{\operatorname{Geom}}(T_{0},T_{1},\vec{x},\vec{y})=\Omega(T_{0},T_{1},x_{1},y_{1})\times\cdots\times\Omega(T_{0},T_{1},x_{k},y_{k}).$
This measure is well-defined provided that $\Omega(T_{0},T_{1},x_{i},y_{i})$ are non-empty for $i\in\llbracket 1,k\rrbracket$, which holds if $y_{i}\geq x_{i}$ for all $i\in\llbracket 1,k\rrbracket$.
The following definition introduces the notion of an $(f,g)$-interlacing geometric line ensemble, which in simple terms is a collection of $k$ independent geometric bridges, conditioned on interlacing with each other and the graphs of two functions $f$ and $g$.
Definition 2.5.
Let $k\in\mathbb{N}$ and $\mathfrak{W}_{k}$ denote the set of signatures of length $k$, i.e.
(2.1) $\mathfrak{W}_{k}=\{\vec{x}=(x_{1},\dots,x_{k})\in\mathbb{Z}^{k}:x_{1}\geq x_{2}\geq\cdots\geq x_{k}\}.$
Let $\vec{x},\vec{y}\in\mathfrak{W}_{k}$, $T_{0},T_{1}\in\mathbb{Z}$ with $T_{0}<T_{1}$, and $f:\llbracket T_{0},T_{1}\rrbracket\rightarrow(-\infty,\infty]$ and $g:\llbracket T_{0},T_{1}\rrbracket\rightarrow[-\infty,\infty)$ be two functions. With the above data we define the $(f,g)$-interlacing geometric line ensemble on the interval $\llbracket T_{0},T_{1}\rrbracket$ with entrance data $\vec{x}$ and exit data $\vec{y}$ to be the $\Sigma$-indexed geometric line ensemble $\mathfrak{Q}$ with $\Sigma=\llbracket 1,k\rrbracket$ on $\llbracket T_{0},T_{1}\rrbracket$ and with the law of $\mathfrak{Q}$ equal to $\mathbb{P}^{T_{0},T_{1},\vec{x},\vec{y}}_{\operatorname{Geom}}$ (the law of $k$ independent uniform increasing paths $\{B_{i}:\llbracket T_{0},T_{1}\rrbracket\rightarrow\mathbb{Z}\}_{i=1}^{k}$ from $B_{i}(T_{0})=x_{i}$ to $B_{i}(T_{1})=y_{i}$), conditioned on
(2.2) $\begin{split}\mathsf{Inter}=&\left\{B_{i}(r-1)\geq B_{i+1}(r)\mbox{ for all $r\in\llbracket T_{0}+1,T_{1}\rrbracket$ and $i\in\llbracket 0,k\rrbracket$}\right\},\end{split}$
with the convention that $B_{0}(x)=f(x)$ and $B_{k+1}(x)=g(x)$.
The above definition is well-posed if there exist $B_{i}\in\Omega(T_{0},T_{1},x_{i},y_{i})$ for $i\in\llbracket 1,k\rrbracket$ that satisfy the conditions in $\mathsf{Inter}$. We denote by $\Omega_{\mathsf{Inter}}(T_{0},T_{1},\vec{x},\vec{y},f,g)$ the set of collections of $k$ increasing paths that satisfy the conditions in $\mathsf{Inter}$ and then the distribution of $\mathfrak{Q}$ is simply the uniform measure on $\Omega_{\mathsf{Inter}}(T_{0},T_{1},\vec{x},\vec{y},f,g)$. We denote the probability distribution of $\mathfrak{Q}$ as $\mathbb{P}_{\mathsf{Inter},\operatorname{Geom}}^{T_{0},T_{1},\vec{x},\vec{y},f,g}$ and write $\mathbb{E}_{\mathsf{Inter},\operatorname{Geom}}^{T_{0},T_{1},\vec{x},\vec{y},f,g}$ for the expectation with respect to this measure. If $f=+\infty$ and $g=-\infty$, we drop them from the notation and simply write $\Omega_{\mathsf{Inter}}(T_{0},T_{1},\vec{x},\vec{y})$, $\mathbb{P}^{T_{0},T_{1},\vec{x},\vec{y}}_{\mathsf{Inter},\operatorname{Geom}}$, and $\mathbb{E}^{T_{0},T_{1},\vec{x},\vec{y}}_{\mathsf{Inter},\operatorname{Geom}}$.
The following definition introduces the notion of the interlacing Gibbs property.
Definition 2.6.
Fix a set $\Sigma=\llbracket 1,N\rrbracket$ with $N\in\mathbb{N}$ or $N=\infty$ and $T_{0},T_{1}\in\mathbb{Z}$ with $T_{0}\leq T_{1}$. A $\Sigma$-indexed geometric line ensemble $\mathfrak{L}:\Sigma\times\llbracket T_{0},T_{1}\rrbracket\rightarrow\mathbb{Z}$ is said to satisfy the interlacing Gibbs property if it is interlacing, meaning that
$L_{i}(j-1)\geq L_{i+1}(j)\mbox{ for all $i\in\llbracket 1,N-1\rrbracket$ and $j\in\llbracket T_{0}+1,T_{1}\rrbracket$},$
and for any finite $K=\llbracket k_{1},k_{2}\rrbracket\subset\llbracket 1,N-1\rrbracket$ and $a,b\in\llbracket T_{0},T_{1}\rrbracket$ with $a<b$ the following holds. Suppose that $f,g$ are two increasing paths drawn in $\{(r,z)\in\mathbb{Z}^{2}:a\leq r\leq b\}$ and $\vec{x},\vec{y}\in\mathfrak{W}_{k}$ with $k=k_{2}-k_{1}+1$ altogether satisfy that $\mathbb{P}(A)>0$ where $A$ denotes the event
$A=\{\vec{x}=({L}_{k_{1}}(a),\dots,{L}_{k_{2}}(a)),\vec{y}=({L}_{k_{1}}(b),\dots,{L}_{k_{2}}(b)),L_{k_{1}-1}\llbracket a,b\rrbracket=f,L_{k_{2}+1}\llbracket a,b\rrbracket=g\},$
where if $k_{1}=1$ we adopt the convention $f=\infty=L_{0}$. Then, writing $k=k_{2}-k_{1}+1$, we have for any $\{B_{i}\in\Omega(a,b,x_{i},y_{i})\}_{i=1}^{k}$ that
(2.3) $\mathbb{P}\left(L_{i+k_{1}-1}\llbracket a,b\rrbracket=B_{i}\mbox{ for $i\in\llbracket 1,k\rrbracket$}\,|A\,\right)=\mathbb{P}_{\mathsf{Inter},\operatorname{Geom}}^{a,b,\vec{x},\vec{y},f,g}\left(\cap_{i=1}^{k}\{Q_{i}=B_{i}\}\right).$
Remark 2.7.
In simple words, a geometric line ensemble is said to satisfy the interlacing Gibbs property if the distribution of any finite number of consecutive paths, conditioned on their end-points and the paths above and below them is simply the uniform measure on all collections of increasing paths that have the same end-points and interlace with each other and the two paths above and below.
Remark 2.8.
Observe that in Definition 2.6 the index $k_{2}$ is assumed to be less than or equal to $N-1$, so that if $N<\infty$ the $N$-th path is special and is not conditionally uniform. If $N=1$, then the conditions in Definition 2.6 become void, i.e., any geometric line ensemble with one line satisfies the geometric Gibbs property. We mention that since $\mathfrak{L}$ is interlacing we automatically have for each $j\in\llbracket T_{0},T_{1}\rrbracket$ that $L_{1}(j)\geq L_{2}(j)\geq\cdots$. In addition, the well-posedness of $\mathbb{P}_{\mathsf{Inter},\operatorname{Geom}}^{T_{0},T_{1},\vec{x},\vec{y},f,g}$ in (2.3) is a consequence of our assumption that $\mathbb{P}(A)>0$.
Remark 2.9.
An immediate consequence of Definition 2.6 is that if $M\leq N$, we have that the induced law on $\{L_{i}\}_{i=1}^{M}$ also satisfies the interlacing Gibbs property as an $\llbracket 1,M\rrbracket$-indexed geometric line ensemble on $\llbracket T_{0},T_{1}\rrbracket$.
2.2. Some properties
In this section we state some of the properties of line ensembles with law $\mathbb{P}_{\mathsf{Inter},\operatorname{Geom}}^{T_{0},T_{1},\vec{x},\vec{y},f,g}$ as in Definition 2.5. The first result is a lemma that shows that these ensembles can be monotonically coupled in their boundary data. Its proof is given in Appendix A.1.
Lemma 2.10.
Assume the same notation as in Definition 2.5. Fix $k\in\mathbb{N}$, $T_{0},T_{1}\in\mathbb{Z}$ with $T_{0}<T_{1}$, two functions $g^{t},g^{b}:\llbracket T_{0},T_{1}\rrbracket\rightarrow[-\infty,\infty)$ as well as $\vec{x}\,^{b},\vec{y}\,^{b},\vec{x}\,^{t},\vec{y}\,^{t}\in\mathfrak{W}_{k}$. We assume that
$g^{t}(r)\geq g^{b}(r)\mbox{ for }r\in\llbracket T_{0},T_{1}\rrbracket\mbox{, }x_{i}^{t}\geq x^{b}_{i}\mbox{, and }y_{i}^{t}\geq y_{i}^{b}\mbox{ for }i=1,\dots,k,$
and also that $\Omega_{\mathsf{Inter}}(T_{0},T_{1},\vec{x}\,^{b},\vec{y}\,^{b},\infty,g^{b})$ and $\Omega_{\mathsf{Inter}}(T_{0},T_{1},\vec{x}\,^{t},\vec{y}\,^{t},\infty,g^{t})$ are both non-empty. Then, there exists a probability space $(\Omega,\mathcal{F},\mathbb{P})$, which supports two $\llbracket 1,k\rrbracket$-indexed geometric line ensembles $\mathfrak{Q}^{t}$ and $\mathfrak{Q}^{b}$ on $\llbracket T_{0},T_{1}\rrbracket$ such that the law of $\mathfrak{Q}^{t}$ (resp. $\mathfrak{Q}^{b}$) under $\mathbb{P}$ is given by $\mathbb{P}_{\mathsf{Inter},\operatorname{Geom}}^{T_{0},T_{1},\vec{x}\,^{t},\vec{y}\,^{t},\infty,g^{t}}$ (resp. $\mathbb{P}_{\mathsf{Inter},\operatorname{Geom}}^{T_{0},T_{1},\vec{x}\,^{b},\vec{y}\,^{b},\infty,g^{b}}$) and such that $\mathbb{P}$-almost surely $Q_{i}^{t}(r)\geq Q^{b}_{i}(r)$ for all $i\in\llbracket 1,k\rrbracket$ and $r\in\llbracket T_{0},T_{1}\rrbracket$.
We next state a strong coupling between geometric random walk bridges and Brownian bridges, which follows from the results in [DW21a]. If $W_{t}$ denotes a standard one-dimensional Brownian motion and $\sigma>0$, then the process
(2.4) $B^{\sigma}_{t}=\sigma(W_{t}-tW_{1}),\hskip 14.22636pt0\leq t\leq 1,$
is called a Brownian bridge (from $B^{\sigma}_{0}=0$ to $B_{1}^{\sigma}=0$) with variance $\sigma^{2}$. When $\sigma^{2}=1$ we drop it from the notation in (2.4) and refer to the latter process as a standard Brownian bridge.
With the above notation we state the strong coupling result we use.
Proposition 2.11.
Fix $p\in(0,\infty)$ and set $\sigma=\sqrt{p(p+1)}$. For any $\epsilon,A>0$ we can find $N_{0}$ large enough, depending on $p,\epsilon,A$, such that the following holds. For any $n\geq N_{0}$ and $z\in\mathbb{Z}_{\geq 0}$ such that $|z-pn|\leq A\sqrt{n}$ we can find a probability space that supports a Brownian bridge $B^{\sigma}$ with variance $\sigma^{2}$ and a random increasing path $\ell^{(n,z)}$ with law $\mathbb{P}_{\operatorname{Geom}}^{0,n,0,z}$ such that
[…]